# Decomposing the Automorphism group of the surreal numbers

### v1

@@ 1. Introduction
where $\alpha$ is an ordinal number, $(y_{\beta})_{\beta<\alpha}$ is a decreasing sequence in $\mathbf{No}$ and $r_{\beta}$ are non-zero coefficients from $\mathbb{R}$ (see [9, Theorem 21], [14, Theorem 5.6], [13, Theorem 14]). Via the Conway normal form, the field of surreal numbers can be regarded as a field of generalized power series – a concept originally dating back to Hahn’s influential work [15] on non-archimedean ordered systems. Even Gonshor himself points out that “for many purposes we can simply work with these generalized power series and ignore what surreal numbers are in the first place” [14, page 71]. The largest term $r_{0}\cdot\omega^{y_{0}}$ in the Conway normal form of $a\neq 0$ determines both the archimedean equivalence class and the sign of $a$: the element $\omega^{y_{0}}$ is the simplest positive representative of the archimedean equivalence class of $a$, and $a$ has positive sign if and only if $r_{0}$ is positive. From a valuation theoretic point of view, this gives rise to $\mathbf{No}$ as the ordered Hahn field $\mathbb{R}\!\left(\!\left(\mathbf{No}\right)\!\right)$ with ordered residue field $\mathbb{R}$ and ordered additive value group $\mathbf{No}$ (under the natural valuation).
In this work, we initiate a study of the automorphism group of the field $\mathbf{No}$. From an algebraic perspective, the motivation behind studying automorphism groups of fields is well-known, in particular in the context of Galois Theory. Especially for a highly complex field such as the surreals, we hope that the systematic analysis of automorphisms sheds further light on the (algebraic) structure and symmetries of $\mathbf{No}$. The more recent works [17] and [18] (based on the third author’s doctoral thesis [20]) prompted us to start our investigation. In these works, the automorphism groups of Hahn fields and Hahn groups are systematically dissected. Based on this structural analysis of automorphism groups, we derive our main result (Theorem 3.11) presenting a decomposition of the automorphism group of the field $\mathbf{No}$ into the product of the following five factors (where semidirect products are particularly highlighted): [^1] The group operation on automorphism groups is given by composition $\circ$ and on vector spaces by ++.
$\operatorname{Aut}(\mathbf{No},+,\cdot)\simeq[1\text{-}\!\operatorname{Aut}(\mathbf{No},+,\cdot)\rtimes\operatorname{Hom}((\mathbf{No},+),(\mathbb{R},+))]\rtimes\left[\operatorname{Int}(\mathbf{No},+)\rtimes\left[\operatorname{Aut}(\mathbf{No},<)\times\mathbb{R}^{\mathbf{No}}\right]\right].$ (1.1)
@@
Now the last factor $\operatorname{Aut}(\mathbf{No},+,<)$ consists of all automorphisms on $\mathbf{No}$ preserving its ordered additive group structure. Consequently, we regard $\mathbf{No}$ as the ordered additive Hahn group $(\mathbb{R}\!\left(\!\left(\mathbf{No}\right)\!\right),+,<)$ for which we obtain in Proposition 3.9 a decomposition into three factors as follows:
@@
In Section 4, we focus on the first factor of the decomposition, in particular, strongly linear automorphisms (automorphisms that commute with infinite sums). We discuss the connection between strongly linear automorphisms and certain derivations, presented in [3]. We use this correspondence to construct strongly linear automorphisms (Example 4.3). We also construct an automorphism that is not strongly linear (Example 4.4), and we use this to show that strongly linear automorphisms are not dense in the group of all automorphisms with respect to the pointwise convergence topology (Remark 4.6). This partially answers a question of Matusinski. In Section 5 we consider automorphisms that preserve further order and exponential structure on the surreals. We show that automorphisms preserving the total ordering and simplicity ordering are trivial (Proposition 5.1), that field automorphisms preserving the omega-map and valuation are trivial (Proposition 5.5), and that strongly linear 1-automorphisms preserving the exponential map are trivial (Proposition 5.2). We leave as questions a complete description of the automorphisms of these richer structures.
@@ 2.1. General conventions
All orderings we consider are linear orderings. Let $(A,<)$ and $(B,<)$ be ordered classes. We use standard interval notation such as $A^{>a}$ for $\{x\in A\mid a<x\}$. A map $f\colon A\to B$ is order preserving if for any $x,y\in A$ with $x<y$ also $f(x)<f(y)$. Given two additive abelian groups $(C,+)$ and $(D,+)$, we denote by $\operatorname{Hom}((C,+),(D,+))$ the group of homomorphism from $C$ to $D$, where the group operation on $\operatorname{Hom}((C,+),(D,+))$ is given by pointwise addition. More generally, given a vector space $V$, we always consider it as the group $(V,+)$ with addition as the group operation. Given a field $k$, we let $k^{\times}\coloneqq k\setminus\{0\}$ denote the group of invertible elements of $k$.
@@ 2.3. Automorphisms
In the rest of this paragraph, we define some groups of automorphisms that will be relevant in the sequel. We refer to [17, Section 3] and [20, Chapter 2] for details.
@@
for some $\rho\in\operatorname{Aut}(k,+,\cdot)$ and $\tau\in\operatorname{Aut}(G,+,<)$. The group $\operatorname{Int}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ of internal (field) automorphisms consists of all $\sigma\in v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ inducing the identity maps, i.e. $\sigma_{k}=\operatorname{id}_{k}$ and $\sigma_{G}=\operatorname{id}_{G}$. Its subgroup $1\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ of $1$-automorphisms consists of all $\sigma\in\operatorname{Int}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ with $\sigma(a)]_{v(a)}=a_{v(a)}$ for any $a\in k\!\left(\!\left(G\right)\!\right)$.
@@
An automorphism $\sigma\in v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ is internal if and only if for any $s\in k\!\left(\!\left(G\right)\!\right)$ the following holds: $v(\sigma(s))=v(s)$, and if $v(s)=0$, then $\sigma(s)]_{0}=s_{0}$. Thus, an automorphism $\sigma\in v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ is a $1$-automorphism if and only if for any $s\in k\!\left(\!\left(G\right)\!\right)$ we have $s_{v(s)}t^{v(s)}=\sigma(s)_{v(\sigma(s))}t^{v(\sigma(s))}$. Alternatively, $\sigma$ is a $1$-automorphism if and only if $\sigma$ fixes the leading term of each $s\in k\!\left(\!\left(G\right)\!\right)$.
@@
An automorphism $\tau\in v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+)$ is internal if and only if for any $s\in k\!\left(\!\left(G\right)\!\right)$ we have $s_{v(s)}t^{v(s)}=[\tau(s)]_{v(\tau(s))}t^{v(\tau(s))}$. Thus, $1\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)\leqslant\operatorname{Int}(k\!\left(\!\left(G\right)\!\right),+)$, and indeed,
@@ 3. Decomposing the automorphism group
Suppose that $k\!\left(\!\left(G\right)\!\right)$ is real closed. Then $\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)=v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$. In particular,
@@
Let $\chi\in\operatorname{Hom}((\mathbf{No},+),(\mathbb{R}^{\times},\cdot))$. Then for any $a\in\mathbf{No}^{\times}$ we have $\chi(a)=\chi(2\cdot a/2)=\chi(a/2)^{2}\in\mathbb{R}^{>0}$. We obtain $\operatorname{Hom}((\mathbf{No},+),(\mathbb{R}^{\times},\cdot))=\operatorname{Hom}((\mathbf{No},+),(\mathbb{R}^{>0},\cdot))$. The lemma follows immediately from the fact that the standard exponential map $\exp$ on $\mathbb{R}$ is an order preserving isomorphism from $(\mathbb{R},+)$ to $(\mathbb{R}^{>0},\cdot)$. ∎
@@
Proposition 3.4 and Corollary 3.6 present the first step of the structural decomposition of $\operatorname{Aut}(\mathbf{No},+,\cdot)$ into three (non-trivial) factors. Indeed, examples of $1$-automorphisms are given in Example 4.3 below, we give a full description of $\operatorname{Hom}((\mathbf{No},+),(\mathbb{R},+))$ in Remark 3.8, and we further decompose $\operatorname{Aut}(\mathbf{No},+,<)$ in Proposition 3.9. First, let us demonstrate exactly how to construct an element of $\operatorname{Aut}(\mathbf{No},+,\cdot)$ from members of these three groups.
@@
Let $\sigma\in 1\text{-}\!\operatorname{Aut}(\mathbf{No},+,\cdot)$, $\varphi\in\operatorname{Hom}((\mathbf{No},+),(\mathbb{R},+))$ and $\tau\in\operatorname{Aut}(\mathbf{No},+,<)$. Then we obtain an automorphism $\theta(\sigma,\varphi,\tau)$ corresponding to the decomposition in Corollary 3.6 as follows:
@@
(see [20, Theorem 2.2.17] and [18, Theorem 3.15]). Moreover, the isomorphism $\Xi\colon\operatorname{Int}(\mathbf{No},+)\rtimes\mathrm{Ext}(\mathbf{No},+)\to v\text{-}\!\operatorname{Aut}(\mathbf{No},+)$ is given by $(\tau_{1},\tau_{2})\mapsto\tau_{1}\circ\tau_{2}$. Now $\operatorname{Int}(\mathbf{No},+)\leqslant\operatorname{Aut}(\mathbf{No},+,<)\leqslant v\text{-}\!\operatorname{Aut}(\mathbf{No},+)$ by Remark 2.8 and [18, Remark 2.11], respectively. Hence,
@@
Finally, note that $\operatorname{Aut}(\mathbb{R},+,<)$ consists of all maps $\mathbb{R}\to\mathbb{R},x\mapsto ax$ for some $a\in\mathbb{R}^{>0}$. Hence, $\operatorname{Aut}(\mathbb{R},+,<)\simeq(\mathbb{R}^{>0},\cdot)$. As $(\mathbb{R}^{>0},\cdot)$ is isomorphic to $(\mathbb{R},+)$ via the real logarithm map $\log$, we obtain
@@
Recall that in 3.7 we presented how $\theta(\sigma,\varphi,\tau)\in\operatorname{Aut}(\mathbf{No},+,\cdot)$ arises from some $\sigma\in 1\text{-}\!\operatorname{Aut}(\mathbf{No},\allowbreak+,\allowbreak\cdot)$, $\varphi\in\operatorname{Hom}((\mathbf{No},+),(\mathbb{R},+))$ and $\tau\in\operatorname{Aut}(\mathbf{No},+,<)$. Based on the decomposition of $\operatorname{Aut}(\mathbf{No},+,<)$ in Proposition 3.9, we refine $\tau$ in the following construction as $\tau(\nu,\zeta,g)$ for some internal group automorphism $\nu\in\operatorname{Int}(\mathbf{No},+)$, some order preserving bijection $\zeta\in\operatorname{Aut}(\mathbf{No},<)$ and some map $g\colon\mathbf{No}\to\mathbb{R}$.
@@ 4. Strongly additive automorphisms
Fact 4.2 ([3, Theorem 3.17]).
If $\partial$ a contracting strongly $k$-linear derivation, then the map $\exp(\partial)\colon k\!\left(\!\left(G\right)\!\right)\to k\!\left(\!\left(G\right)\!\right)$ given by
@@
In many cases, it is easier to work with contractive strongly linear derivations than $1$-automorphisms, as we will see in Proposition 5.2 below. It is also easy to find nontrivial contractive derivations, using the following variant of a construction from [8, 16]:
@@
Then $\partial_{\varphi}$ is a contractive strongly $\mathbb{R}$-linear derivation, so $\exp(\partial_{\varphi})$ belongs to $1\text{-}\!\operatorname{Aut}(\mathbf{No},+,\cdot)$.
We end this section by giving an example of a $1$-automorphism of $(\mathbf{No},+,\cdot)$ that is not strongly additive:
@@
Using global choice, fix a well-ordered $\mathbb{Q}$-vector space basis $(y_{\alpha})_{\alpha\in\mathbf{On}}$ of $\mathbf{No}$ with $y_{0}=1$ and $y_{1}=\omega$. For each $\alpha\in\mathbf{On}$, let $\Gamma_{\alpha}\coloneqq\bigoplus_{\beta<\alpha}\mathbb{Q}y_{\beta}$, and let $K_{\alpha}\coloneqq\mathbb{R}\!\left(\!\left(\Gamma_{\alpha}\right)\!\right)$, identified in the usual way with a subfield of $\mathbf{No}$. We define an automorphism $\sigma_{\alpha}$ on $K_{\alpha}$ for each $\alpha\geqslant 2$ as follows:
@@
    The field $K_{2}=K\!\left(\!\left(\mathbb{Q}\oplus\mathbb{Q}\omega\right)\!\right)$ is an immediate extension of $K(a,\omega^{\omega})^{\operatorname{rc}}=K(a,\varepsilon)^{\operatorname{rc}}$. Using that $\omega^{\omega}$ is greater than both $K(a)^{\operatorname{rc}}$ and $K(a+\varepsilon)^{\operatorname{rc}}$, we extend $\sigma_{1}$ to an automorphism of $K(a,\omega^{\omega})^{\operatorname{rc}}$ that fixes $\omega^{\omega}$. We further extend this to an automorphism $\sigma_{2}$ of $K_{2}$ using the uniqueness of immediate extensions.
@@
    Let $\alpha\geqslant 2$ and assume that we have defined $\sigma_{\alpha}$. Then $K_{\alpha+1}$ is an immediate extension of $K_{\alpha}(\omega^{y_{\alpha}})^{\operatorname{rc}}$. We first extend $\sigma_{\alpha}$ to an automorphism of $K_{\alpha}(\omega^{y_{\alpha}})^{\operatorname{rc}}$ that fixes $\omega^{y_{\alpha}}$, and then we further extend this to $K_{\alpha+1}$ by uniqueness of immediate extensions.
@@
    Let $\lambda>0$ be a limit ordinal, and assume we’ve defined $\sigma_{\alpha}$ for $\alpha<\lambda$. Then $K_{\lambda}$ is an immediate extension of $\bigcup_{\alpha<\lambda}K_{\alpha}$, so the automorphisms $(\sigma_{\alpha})_{\alpha<\lambda}$ extend to an automorphism $\sigma_{\lambda}$ of $K_{\lambda}$, again using uniqueness of immediate extensions.
@@
Strongly linear automorphisms have the desirable property that they can be easily expressed by declaring them on monomials and extending by strong linearity. It is thus natural to ask whether non-strongly linear automorphisms can be approximated by strongly linear ones. The following is a special case of a question asked by M. Matusinski for the general case of Hahn fields.
@@
$U_{\sigma}(a,\omega)=\{\tau\in\operatorname{Aut}(\mathbf{No},+,\cdot):\tau(a)=a+\omega^{-\omega}\text{ and }\tau(\omega)=\omega\}.$

### v2

@@ 1. Introduction
where $\alpha$ is an ordinal number, $(y_{\beta})_{\beta<\alpha}$ is a strictly decreasing sequence in $\mathbf{No}$ and $r_{\beta}$ are non-zero coefficients from $\mathbb{R}$ (see [9, Theorem 21], [14, Theorem 5.6], [13, Theorem 14]). Via the Conway normal form, the field of surreal numbers can be regarded as a field of generalized power series—a concept originally dating back to Hahn’s influential work [15] on non-archimedean ordered systems. Even Gonshor himself points out that “for many purposes we can simply work with these generalized power series and ignore what surreal numbers are in the first place” [14, page 71]. The largest term $r_{0}\cdot\omega^{y_{0}}$ in the Conway normal form of $a\neq 0$ determines both the archimedean equivalence class and the sign of $a$: the element $\omega^{y_{0}}$ is the simplest positive representative of the archimedean equivalence class of $a$, and $a$ has positive sign if and only if $r_{0}$ is positive. From a valuation theoretic point of view, this gives rise to $\mathbf{No}$ as the ordered Hahn field $\mathbb{R}\!\left(\!\left(\mathbf{No}\right)\!\right)$ with ordered residue field $\mathbb{R}$ and ordered additive value group $\mathbf{No}$ (under the natural valuation).
In this work, we initiate a study of the automorphism group of the field $\mathbf{No}$. From an algebraic perspective, the motivation behind studying automorphism groups of fields is well-known, in particular in the context of Galois Theory. Especially for a highly complex field such as the surreals, we hope that the systematic analysis of automorphisms sheds further light on the (algebraic) structure and symmetries of $\mathbf{No}$. The more recent works [17] and [18] (based on the third author’s doctoral thesis [20]) prompted us to start our investigation. In these works, the automorphism groups of Hahn fields and Hahn groups are systematically dissected. Based on this structural analysis of automorphism groups, we derive our main result (Theorem 3.11) presenting a decomposition of the automorphism group of the field $\mathbf{No}$ into the product of the following five factors (where semidirect products are particularly highlighted): [^1] The group operation on automorphism groups is given by composition $\circ$ and on vector spaces by addition ++.
$\operatorname{Aut}(\mathbf{No},+,\cdot)\simeq[1\text{-}\!\operatorname{Aut}(\mathbf{No},+,\cdot)\rtimes\operatorname{Hom}((\mathbf{No},+),(\mathbb{R},+))]\rtimes\left[\operatorname{Int}(\mathbf{No},+)\rtimes\left[\operatorname{Aut}(\mathbf{No},<)\times\mathbb{R}^{\mathbf{No}}\right]\right]$ (1.1)
@@
Now the last factor $\operatorname{Aut}(\mathbf{No},+,<)$ consists of all automorphisms on $\mathbf{No}$ preserving its ordered additive group structure. Consequently, we regard $\mathbf{No}$ as the ordered additive Hahn group $(\mathbb{R}\!\left(\!\left(\mathbf{No}\right)\!\right),+,<)$, for which we obtain in Proposition 3.9 a decomposition into three factors as follows:
@@
In Section 4, we focus on automorphisms in the first factor of decomposition (1.1), in particular, strongly linear ones (automorphisms that commute with infinite sums). We discuss the connection between strongly linear automorphisms and certain derivations on $\mathbf{No}$, based on the correspondence presented in [6]. This correspondence enables us to construct strongly linear automorphisms (Example 4.3). We also construct an automorphism that is not strongly linear (Example 4.4), and we use this to show that strongly linear automorphisms are not dense in the group of all automorphisms with respect to the pointwise convergence topology (Remark 4.6). In Section 5, we consider automorphisms that preserve further order and exponential structure on the surreals. We show that automorphisms preserving the total ordering and simplicity ordering are trivial (Proposition 5.1), that field automorphisms preserving the omega-map and valuation are trivial (Proposition 5.5), and that strongly linear 1-automorphisms preserving the exponential map are trivial (Proposition 5.2). We leave as questions a complete description of the automorphisms of these richer structures.
@@ 2.1. General conventions
All orderings we consider are linear orderings, the one exception being in Proposition 5.1. Let $(A,<)$ and $(B,<)$ be ordered classes. We use standard interval notation such as $A^{>a}$ for $\{x\in A\mid a<x\}$. A map $f\colon A\to B$ is order preserving if for any $x,y\in A$ with $x<y$ also $f(x)<f(y)$. Given two additive abelian groups $(C,+)$ and $(D,+)$, we denote by $\operatorname{Hom}((C,+),(D,+))$ the group of homomorphism from $C$ to $D$, where the group operation on $\operatorname{Hom}((C,+),(D,+))$ is given by pointwise addition. More generally, given a vector space $V$, we always consider it as the group $(V,+)$ with addition as the group operation. Given a field $k$, we let $k^{\times}\coloneqq k\setminus\{0\}$ denote the group of invertible elements of $k$.
@@ 2.3. Automorphisms
In the rest of this subsection, we define some groups of automorphisms that will be relevant in the sequel. We refer to [17, Section 3] and [20, Chapter 2] for details.
@@
for some $\rho\in\operatorname{Aut}(k,+,\cdot)$ and $\tau\in\operatorname{Aut}(G,+,<)$. The group $\operatorname{Int}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ of internal (field) automorphisms consists of all $\sigma\in v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ inducing the identity maps, i.e. $\sigma_{k}=\operatorname{id}_{k}$ and $\sigma_{G}=\operatorname{id}_{G}$. Its subgroup $1\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ of $1$-automorphisms consists of all $\sigma\in\operatorname{Int}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ with $\sigma(a)]_{v(a)}=a_{v(a)}$ for any $a\in k\!\left(\!\left(G\right)\!\right)^{\times}$.
@@
An automorphism $\sigma\in v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ is internal if and only if for any $s\in k\!\left(\!\left(G\right)\!\right)$ the following holds: $v(\sigma(s))=v(s)$, and if $v(s)=0$, then $\sigma(s)]_{0}=s_{0}$. Thus, an automorphism $\sigma\in v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$ is a $1$-automorphism if and only if for any $s\in k\!\left(\!\left(G\right)\!\right)^{\times}$ we have $s_{v(s)}t^{v(s)}=\sigma(s)_{v(\sigma(s))}t^{v(\sigma(s))}$. Alternatively, $\sigma$ is a $1$-automorphism if and only if $\sigma$ fixes the leading term of each $s\in k\!\left(\!\left(G\right)\!\right)^{\times}$.
@@
An automorphism $\tau\in v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+)$ is internal if and only if for any $s\in k\!\left(\!\left(G\right)\!\right)^{\times}$ we have $s_{v(s)}t^{v(s)}=[\tau(s)]_{v(\tau(s))}t^{v(\tau(s))}$. Thus, $1\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)\leqslant\operatorname{Int}(k\!\left(\!\left(G\right)\!\right),+)$, and indeed,
@@ 3. Decomposing the automorphism group
Suppose that $k\!\left(\!\left(G\right)\!\right)$ is real closed and that $k$ is archimedean. Then $\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)=v\text{-}\!\operatorname{Aut}(k\!\left(\!\left(G\right)\!\right),+,\cdot)$. In particular,
@@
Let $\chi\in\operatorname{Hom}((\mathbf{No},+),(\mathbb{R}^{\times},\cdot))$. Then for any $a\in\mathbf{No}$ we have $\chi(a)=\chi(2\cdot a/2)=\chi(a/2)^{2}\in\mathbb{R}^{>0}$. We obtain $\operatorname{Hom}((\mathbf{No},+),(\mathbb{R}^{\times},\cdot))=\operatorname{Hom}((\mathbf{No},+),(\mathbb{R}^{>0},\cdot))$. The lemma follows immediately from the fact that the standard exponential map $\exp$ on $\mathbb{R}$ is an order preserving isomorphism from $(\mathbb{R},+)$ to $(\mathbb{R}^{>0},\cdot)$. ∎
@@
Proposition 3.4 and Corollary 3.6 present the first step of the structural decomposition of $\operatorname{Aut}(\mathbf{No},+,\cdot)$ into three (non-trivial) factors. Indeed, examples of non-trivial $1$-automorphisms are given in Example 4.3, we give a full description of $\operatorname{Hom}((\mathbf{No},+),(\mathbb{R},+))$ in Remark 3.8, and we further decompose $\operatorname{Aut}(\mathbf{No},+,<)$ in Proposition 3.9. First, let us demonstrate exactly how to construct an element of $\operatorname{Aut}(\mathbf{No},+,\cdot)$ from members of these three groups.
@@
Let $\sigma\in 1\text{-}\!\operatorname{Aut}(\mathbf{No},+,\cdot)$, $\varphi\in\operatorname{Hom}((\mathbf{No},+),(\mathbb{R},+))$ and $\tau\in\operatorname{Aut}(\mathbf{No},+,<)$. Then we obtain an automorphism $\theta(\sigma,\varphi,\tau)$ corresponding to the decomposition in Corollary 3.6 as follows.
@@
(see [20, Theorem 2.2.17] and [18, Theorem 3.15]). Moreover, the isomorphism $\Xi\colon\operatorname{Int}(\mathbf{No},+)\rtimes\mathrm{Ext}(\mathbf{No},+)\to v\text{-}\!\operatorname{Aut}(\mathbf{No},+)$ is given by $(\tau_{1},\tau_{2})\mapsto\tau_{1}\circ\tau_{2}$. Now $\operatorname{Int}(\mathbf{No},+)\leqslant\operatorname{Aut}(\mathbf{No},+,<)\leqslant v\text{-}\!\operatorname{Aut}(\mathbf{No},+)$ by Remark 2.8 and Lemma 3.2, respectively. Hence,
@@
Finally, note that the group $\operatorname{Aut}(\mathbb{R},+,<)$ consists of all maps $\mathbb{R}\to\mathbb{R},x\mapsto ax$ for some $a\in\mathbb{R}^{>0}$. Hence, $\operatorname{Aut}(\mathbb{R},+,<)\simeq(\mathbb{R}^{>0},\cdot)$. As $(\mathbb{R}^{>0},\cdot)$ is isomorphic to $(\mathbb{R},+)$ via the real logarithm map $\log$, we obtain
@@
Recall from 3.7 how $\theta(\sigma,\varphi,\tau)\in\operatorname{Aut}(\mathbf{No},+,\cdot)$ arises from some $\sigma\in 1\text{-}\!\operatorname{Aut}(\mathbf{No},\allowbreak+,\allowbreak\cdot)$, $\varphi\in\operatorname{Hom}((\mathbf{No},+),(\mathbb{R},+))$ and $\tau\in\operatorname{Aut}(\mathbf{No},+,<)$. Based on the decomposition of $\operatorname{Aut}(\mathbf{No},+,<)$ in Proposition 3.9, we refine $\tau$ in the following construction as $\tau(\nu,\zeta,g)$ for some internal group automorphism $\nu\in\operatorname{Int}(\mathbf{No},+)$, some order preserving bijection $\zeta\in\operatorname{Aut}(\mathbf{No},<)$ and some map $g\colon\mathbf{No}\to\mathbb{R}$.
@@ 4. Strongly additive automorphisms
Fact 4.2 ([6, Theorem 3.13]).
If $\partial$ is a contracting strongly $k$-linear derivation, then the map $\exp(\partial)\colon k\!\left(\!\left(G\right)\!\right)\to k\!\left(\!\left(G\right)\!\right)$ given by
@@
In many cases, it is easier to work with contracting strongly linear derivations than $1$-automorphisms, as we will see in Proposition 5.2 below. It is also easy to find non-trivial contracting derivations, using the following variant of a construction from [8, 16].
@@
Then $\partial_{\varphi}$ is a contracting strongly $\mathbb{R}$-linear derivation, so $\exp(\partial_{\varphi})$ belongs to $1\text{-}\!\operatorname{Aut}(\mathbf{No},+,\cdot)$.
We now present an example of a $1$-automorphism of $(\mathbf{No},+,\cdot)$ that is not strongly additive.
@@
Using global choice, fix a well-indexed $\mathbb{Q}$-vector space basis $(y_{\alpha})_{\alpha\in\mathbf{On}}$ of $\mathbf{No}$ with $y_{0}=1$ and $y_{1}=\omega$. For each $\alpha\in\mathbf{On}$, let $\Gamma_{\alpha}\coloneqq\bigoplus_{\beta<\alpha}\mathbb{Q}y_{\beta}$, and let $K_{\alpha}\coloneqq\mathbb{R}\!\left(\!\left(\Gamma_{\alpha}\right)\!\right)$, identified in the usual way with a subfield of $\mathbf{No}$. We define an automorphism $\sigma_{\alpha}$ on $K_{\alpha}$ for each $\alpha\geqslant 2$ as follows:
@@
    The field $K_{2}=K\!\left(\!\left(\mathbb{Q}\oplus\mathbb{Q}\omega\right)\!\right)$ is an immediate extension of $K(a,\omega^{\omega})^{\operatorname{rc}}=K(a,\varepsilon)^{\operatorname{rc}}$. Using that $\omega^{\omega}$ is greater than both $K(a)^{\operatorname{rc}}$ and $K(a+\varepsilon)^{\operatorname{rc}}$, we extend $\sigma_{1}$ to an automorphism of $K(a,\omega^{\omega})^{\operatorname{rc}}$ that fixes $\omega^{\omega}$. We further extend this to an automorphism $\sigma_{2}$ of $K_{2}$ using the uniqueness of maximal immediate extensions.
@@
    Let $\alpha\geqslant 2$ and assume that we have defined $\sigma_{\alpha}$. Then $K_{\alpha+1}$ is an immediate extension of $K_{\alpha}(\omega^{y_{\alpha}})^{\operatorname{rc}}$. We first extend $\sigma_{\alpha}$ to an automorphism of $K_{\alpha}(\omega^{y_{\alpha}})^{\operatorname{rc}}$ that fixes $\omega^{y_{\alpha}}$, and then we further extend this to $K_{\alpha+1}$ by uniqueness of maximal immediate extensions.
@@
    Let $\lambda>0$ be a limit ordinal, and assume we have defined $\sigma_{\alpha}$ for $\alpha<\lambda$. Then $K_{\lambda}$ is an immediate extension of $\bigcup_{\alpha<\lambda}K_{\alpha}$, so the automorphisms $(\sigma_{\alpha})_{\alpha<\lambda}$ extend to an automorphism $\sigma_{\lambda}$ of $K_{\lambda}$, again using uniqueness of maximal immediate extensions.
@@
Strongly linear automorphisms have the desirable property that they can be easily expressed by declaring them on monomials and extending by strong linearity. It is thus natural to ask whether non-strongly linear automorphisms can be approximated by strongly linear ones.
@@
$U_{\sigma}(a,\omega)=\{\tau\in\operatorname{Aut}(\mathbf{No},+,\cdot)\mid\tau(a)=a+\omega^{-\omega}\text{ and }\tau(\omega)=\omega\}.$